OpenStudy (prof_cutederp):

im so confused o_O Solve for a: 6 = 9/10a - 1

MsBrains (ms-brains):

I think Cymath.com might be able to help you. :)

OpenStudy (daniel.ohearn1):

Is the problem 6= (9/10)a -1 ?

OpenStudy (prof_cutederp):

yes ;-;

OpenStudy (daniel.ohearn1):

or is it indeed 6=9/(10a) -1

OpenStudy (prof_cutederp):

Solve for a: 6 = 9/10a - 1 just like this

OpenStudy (daniel.ohearn1):

9/10 of a should be written (9/10)a because its different from (9/10a)

OpenStudy (prof_cutederp):

its how it was written im sorry

OpenStudy (daniel.ohearn1):

9 divided by 10a is what the second case would read

OpenStudy (daniel.ohearn1):

Okay lets say the problem is 9/10 of a plus 1 = 6 okay?

OpenStudy (prof_cutederp):

ok ;-;

OpenStudy (daniel.ohearn1):

In algebra we can change each side of the equation so that we are able to solve for a, by manipulating it

OpenStudy (daniel.ohearn1):

First thing we can do is add 1 to each side

OpenStudy (daniel.ohearn1):

So now we have 9/10 of a = 7

OpenStudy (daniel.ohearn1):

Does that make sense to you? I added 1 to both sides and since we subtract 1 and add 1 to the right side that becomes -1 + 1 which equals zero